# Automatic Link Selection in Multi-Channel Multiple Access with Link Failures

### v1

@@ Abstract
This paper focuses on the problem of automatic link selection in multi-channel multiple access control using bandit feedback. In particular, a controller assigns multiple users to multiple channels in a time slotted system, where in each time slot at most one user can be assigned to a given channel and at most one channel can be assigned to a given user. Given that user $i$ is assigned to channel $j$, the transmission fails with a fixed probability $f_{i,j}$. The failure probabilities are not known to the controller. The assignments are made dynamically using success/failure feedback. The goal is to maximize the time average utility, where we consider an arbitrary (possibly nonsmooth) concave and entrywise nondecreasing utility function. The problem of merely maximizing the total throughput has a solution of always assigning the same user-channel pairs and can be unfair to certain users, particularly when the number of channels is less than the number of users. Instead, our scheme allows various types of fairness, such as proportional fairness, maximizing the minimum, or combinations of these by defining the appropriate utility function. We propose two algorithms for this task. The first algorithm is adaptive and gets within $\mathcal{O}(\log(T)/T^{1/3})$ of optimality over any interval of $T$ consecutive slots over which the success probabilities do not change. The second algorithm has faster $\mathcal{O}(\sqrt{\log(T)/T})$ performance over the first $T$ slots, but does not adapt well if probabilities change.
Index Terms:
Multi-armed bandit learning; Proportional fairness; Network utility maximization; Optimization; Stochastic control; Adaptive learning
@@ I Introduction
We consider the Multiple Access Control (MAC) problem with $n$ users and $m$ channels in slotted time $t\in\mathbb{N}$. In each time slot, a controller has to assign the users to channels such that at most one user is assigned to given channel and at most one channel is assigned to a given user. The channel assignments may fail. In particular, there exist $q_{i,j}\in[0,1]$ for each $(i,j)\in\{1,2,\dots,n\}\times\{1,2,\dots,m\}$, where in time slot $t$, given that the controller decided to assign user $i$ to channel $j$, the assignment fails independently with probability $1-q_{i,j}$. The controller does not know the probabilities $q_{i,j}$. Instead, at the end of every slot, it receives feedback on whether the transmission for each assigned user-channel pair succeeded or failed.
@@
$S_{i,j}(t)=\begin{cases}1&\text{ if link }i,j\text{ is successful in time slot }t\\ 0&\text{ otherwise},\end{cases}$
$Y_{i,j}(t)=\begin{cases}1&\text{ user }i\text{ is assigned to channel }j\text{ in time slot }t\\ 0&\text{ otherwise},\end{cases}$
and $X_{i}(t)=\sum_{j=1}^{m}Y_{i,j}(t)S_{i,j}(t)$ for all $i\in\{1,2,\dots,n\}$. The goal is to maximize $\lim_{T\to\infty}\phi(\mathbb{E}\{\boldsymbol{\overline{X}}(T)\})$ using feedback on the link failures, where $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is a concave entrywise nondecreasing utility function known to the controller and $\boldsymbol{\overline{X}}(T)=\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)$.  [^1] The limit is assumed to exist for simplicity of this introduction; the precise goal is to maximize a $\lim\inf_{T\to\infty}\phi(\boldsymbol{\overline{X}}(T))$
@@
In addition, we are looking for algorithms that are adaptive. Formally, consider a system in which the channel success probabilities may change. In such a system, given a $T\in\mathbb{N}$, we require
$\phi^{\text{opt}}-\phi\left(\frac{1}{T}\mathbb{E}\left\{\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)\leq g(T)$
for any $T_{0}\in\mathbb{N}$, irrespective of the success probabilities outside of the time frame $[T_{0}:T_{0}+T-1]$, given that success probabilities remained constant in the frame. Here, $\phi^{\text{opt}}$ is the optimal utility of the original problem that uses the constant success probabilities in $[T_{0}:T_{0}+T-1]$ of the above scenario. Note that $g$ is the same function regardless of $T_{0}$.
@@ I-A Utility functions
Another choice is $\phi(\boldsymbol{x})=w_{1}\min\{x_{1},\dots,x_{n}\}+w_{2}\sum_{i=1}^{n}\log(1+\beta x_{i})$, where $w_{1},w_{2},\beta$ are given nonnegative weights. The logarithmic term introduces a form of proportional fairness [1, 2]. See also discussion of different utility functions in [1, 2, 3, 4, 5]
@@ I-B Related work
Network scheduling in stochastic environments is a widely considered problem in the literature on communication networks. Some examples include scheduling in wireless networks [6, 7], computer networks [8], vehicular networks [9] and unmanned aerial vehicle networks [10]. These works consider different goals such as optimal scheduling for queue-stability [11], power minimization [12], utility maximization [13], ensuring fairness [14], flow control [15], and multiple access control [16].
@@
Multiple access, where many users access a limited number of communication channels, is an important problem in network scheduling. Here, it is desirable for users to be scheduled to avoid collisions. Link failures occur when the receiver is unable to decode the packet transmissions. This can occur, for example, when a fixed transmission power is used, but channel conditions have random and unknown fluctuations so that the received signal strength is insufficient for decoding. Different links can have different properties (such as different geographic distances to the receiver), so they can also have different success probabilities. The naive approach of assigning the users with links with the least failure probabilities leads to unfairness since, in such a scenario, users with high link failure probabilities will never be assigned. A common approach to solve this problem is to maximize a utility function of the time-averaged success. This problem has been well studied in the full information scenario when either the fluctuating channel conditions are known before transmission (called opportunistic scheduling) or when channel success probabilities are known in advance. Opportunistic scheduling has been considered using utility functions [17], Lyapunov drift [18], Frank-Wolfe [19, 20, 21], primal-dual [22, 23], and drift-plus-penalty [24]. The case when success probabilities are known in advance can be solved offline as a convex optimization problem using the mirror descent technique [25].
The problem becomes challenging when the success probabilities are not known, but we only receive bandit feedback on the successes. The problem has to be approached by combining ideas from optimizing functions of time averages with multi-armed bandit learning. The work on bandits with vector rewards and concave utility functions can be adapted for the single channel case ($m=1$) of our problem (See [26, 27, 28]). There are two main drawbacks to these approaches. First, the above works do not consider the matching constraints considered in our work. Next, they focus on upper confidence bound (UCB) techniques and are not adaptive. It is possible to develop adaptive algorithms using techniques from the above works together with the EXP3 algorithm [29] for the single channel case. However, they cannot be directly extended for the multi-channel case. The main reason for this is the complexity of the inner problem arising in each iteration which is a problem over the set of doubly stochastic matrices (Birkhoff polytope). This can be addressed using Sinkhorn’s algorithm [30] in each iteration. However, this results in an algorithm that requires computationally complex inner iterations to be performed within each iteration. The work of [31] uses follow the regularized leader approach to solve adversarial bandit problems over the set of doubly stochastic matrices. However, their algorithm relies on a computationally expensive inner iteration similar to what we described before. The work of  [32] proposes an algorithm to solve online optimization problems over transport polytopes. All the inner iterations of their algorithm have explicit solutions thanks to the rounding trick introduced for transport polytopes by [33]. In our work, we adapt the rounding trick to the Birkhoff polytope and develop an algorithm for general $m$. We also develop a UCB-based algorithm for our problem. It should be noted that although this algorithm achieves faster convergence, it is not adaptive. However, the UCB algorithm has additional advantages such as distributed implementation. Our paper also treats general (possibly nonsmooth) utility functions, so Frank-Wolfe methods that rely on smoothness cannot be used.
@@
Existing work on using learning for scheduling problems in communications include spectrum sharing [34, 35], network utility maximization with unknown utility functions [36], and queue stability [11]. The existing literature on bandit learning-based methods for scheduling problems focuses on maximizing throughput [37]. The work of [36] considers bandit learning with unknown utilities, and the work of [38] considers the online resource allocation problem, where the objective is to maximize the $\alpha$-fairness of the allocations. Neither of the above works considers the bandit reward structure, and both works have a different problem structure from ours. In addition, the above works do not focus on the adaptiveness of the proposed algorithms.
@@
It should be noted that the concept of adaptiveness considered in this paper is different from the adversarial setting [39, 40, 29], although they have similarities. In adversarial settings, optimality is defined with respect to the rewards received throughout the time horizon. This is useful when no stochastic assumptions can be placed on the rewards. On the other hand, the adaptiveness considered in this paper is useful when we can place stochastic assumptions on the rewards, but the distributions may change from time to time. In this case, we can define an optimal strategy for each time frame in which the reward distribution remained constant. Hence, the goal in this frame is to learn the aforementioned strategy irrespective of the reward distributions of the past. An approach commonly used in problems of this flavor is minimizing dynamic regret [41, 42, 43]. Here, the regret is modified to account for the changing environments and the regret bounds are in terms of some measure that captures the degree of change. Various algorithms are developed for the setting with linear utility functions using optimizing in phases/episodes [41, 42], and sliding window-based algorithms [43]. We utilize a simpler notion of adaptiveness and develop an algorithm for the case with general utility functions.
Multi-armed bandit (MAB) learning [44, 45] is a class of extensively studied problems. In the classical MAB setting, a user chooses from multiple arms, each incurring a fixed mean cost. The choice is made without knowledge of the mean costs, using the feedback received on the cost of the chosen arm. The goal is to learn the arm with the lowest mean cost as fast as possible. Typically, the algorithms implemented to solve such problems involve an exploration phase, where the user explores all arms in order to learn the mean costs, after which, in the exploitation phase, the user exploits the learned information to make good decisions [46, 47]. Upper confidence bound-based algorithms, where the algorithm maintains an upper bound on the mean cost of each arm, is a popular approach to solve these problems [47, 48]. Adversarial bandit learning [39, 40], where an adversary is allowed to assign the cost of each arm after the user selects the arm, is an extension of multi-armed bandit learning. In such problems, no stochastic assumptions are placed on the rewards. These problems are solved using algorithms such as the EXP3 algorithm [29], where in each iteration of the algorithm, a probability distribution over the space of all possible actions is found, after which the decision is randomly sampled from the distribution. There are also many extensions of the MAB problem, such as linear bandits, contextual bandits, and combinatorial bandits [47].
@@ I-C Our Contributions
We develop and analyze two algorithms to solve the problem of automatic link selection in multi-channel multiple access, combining the ideas of multi-armed bandit learning and Lyapunov optimization. Although the classical MAB problem has been widely analyzed for maximizing linear utilities, they suffer from lack of fairness in assignments when applied to the considered problem. It is notable that our method allows either smooth or nonsmooth concave, entrywise nondecreasing utilities.
We prove that the first algorithm gets within $\mathcal{O}(T^{-1/3}\log(T))$ of optimality over any interval of $T$ consecutive time slots during which the (unknown) success probabilities do not change. If these probabilities are different before $T_{0}$, but change to new probabilities during $\{T_{0},T_{0}+1,\dots,T_{0}+T-1\}$, our performance guarantees for the new interval are independent of behaviors before $T_{0}$, even though the algorithm does not know the exact time $T_{0}$ of the change. Hence, the algorithm is adaptive. This is possible thanks to the importance sampling technique in [29], which we use to estimate the true link failure probabilities in each time slot. This technique uses feedback only from the previous time slot.
The second algorithm is based on upper confidence bound (UCB) techniques and gets within $\mathcal{O}(\sqrt{\log(T)/T})$ of optimality within a finite time horizon of $T$ time slots. Although this algorithm has faster convergence compared to the first algorithm, it is not adaptive due to the UCB-based estimation. However, the channel-user assignment of this algorithm has a max-weight structure and can be solved using the well-known Hungarian algorithm [49]. This is in contrast to the first algorithm, where the assignment is sampled from a distribution in each time slot, which may be less desirable in certain scenarios. Due to the max-weight decisions, the second algorithm can be implemented in a distributed setting in the absence of a centralized controller, given that all the users have access to feedback on which channels are successfully accessed in each time slot. Simulations also depict the adaptiveness of the first algorithm and the faster convergence of the second algorithm.

### v2

@@ Abstract
This paper focuses on the problem of automatic link selection in multi-channel multiple access control using bandit feedback. In particular, a controller assigns multiple users to multiple channels in a time-slotted system, where in each time slot, at most one user can be assigned to a given channel, and at most one channel can be assigned to a given user. Given that user $i$ is assigned to channel $j$, the transmission fails with a fixed unknown probability $1-q_{i,j}$. The assignments are made dynamically using success/failure feedback. The goal is to maximize the time-average utility, where we consider an arbitrary (possibly nonsmooth) concave, entrywise nondecreasing utility function. The first proposed algorithm has fast $\mathcal{O}(\sqrt{\log(T)/T})$ convergence. However, this algorithm requires solving a convex optimization problem within each iteration, which can be computationally expensive. The second algorithm has slower $\mathcal{O}(\sqrt[3]{\log(T)/T})$ convergence, while avoiding the costly inner optimization. Both of these algorithms are adaptive. In particular, the convergence guarantee holds for any interval of $T$ consecutive slots during which the success probabilities do not change. We further study several special cases. In the single-channel setting, we obtain both fast $\mathcal{O}(\sqrt{\log(T)/T})$ convergence and efficient implementation via a simpler adaptive mechanism. We also consider a UCB-based non-adaptive algorithm with max-weight-type decisions. Simulations highlight intriguing performance trade-offs and demonstrate rapid adaptation of the proposed adaptive schemes.
Automatic Link Selection in Multi-Channel Multiple Access with Link Failures
Mevan Wijewardena, Michael J. Neely, Haipeng Luo Mevan Wijewardena and Michael J. Neely are with the Electrical Engineering department at the University of Southern California. Haipeng Luo is with the Computer Science department at the University of Southern California. This work was presented in part at the WiOpt conference in Linköping, Sweden, May 2025 [50].
@@ I Introduction
We consider the Multiple Access Control (MAC) problem with $n$ users and $m$ channels in slotted time $t\in\mathbb{N}$. In each time slot, a controller assigns users to channels such that at most one user is assigned to a given channel and at most one channel is assigned to a given user. The transmissions over each channel may fail. In particular, there exist $q_{i,j}\in[0,1]$ for each $(i,j)\in\{1,2,\dots,n\}\times\{1,2,\dots,m\}$, where in time slot $t$, given that the controller decided to assign user $i$ to channel $j$, the transmission fails independently with probability $1-q_{i,j}$. The controller does not know the probabilities $q_{i,j}$. Instead, at the end of every slot, it receives feedback on the successes.
@@
[…]
and $X_{i}(t)=\sum_{j=1}^{m}Y_{i,j}(t)S_{i,j}(t)$ for all $i\in\{1,2,\dots,n\}$. The goal is to maximize $\lim_{T\to\infty}\phi(\mathbb{E}\{\boldsymbol{\overline{X}}(T)\})$ using feedback on the transmission failures, where $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is a concave entrywise nondecreasing utility function known to the controller and $\boldsymbol{\overline{X}}(T)=\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)$.  [^1] The limit is assumed to exist for simplicity of this introduction; the precise goal is to maximize a $\lim\inf_{T\to\infty}\phi(\boldsymbol{\overline{X}}(T))$
@@
In addition, we are looking for algorithms that are adaptive. Formally, consider a system in which the channel success probabilities $q_{i,j}$ may change. In such a system, given a $T\in\mathbb{N}$, we require
$\phi^{\text{opt}}-\phi\left(\frac{1}{T}\mathbb{E}\left\{\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)\leq g(T)$ (1)
for any $T_{0}\in\mathbb{N}$, irrespective of the success probabilities outside of the time frame $T_{0}:T_{0}+T-1]$, given that success probabilities remained constant in the frame. Here, $\phi^{\text{opt}}$ is the optimal utility of the original problem that uses the constant success probabilities in $T_{0}:T_{0}+T-1]$ of the above scenario. Note that $g$ is the same function regardless of $T_{0}$. In fact, it can be shown that given $T^{{}^{\prime}}\in\mathbb{N}$, our algorithms can achieve
$\phi^{\text{opt}}-\phi\left(\frac{1}{T}\mathbb{E}\left\{\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)\leq g(T^{{}^{\prime}})$
for any $T_{0}\in\mathbb{N}$ and $T\geq T^{{}^{\prime}}$, where the algorithm parameters only depend on $T^{{}^{\prime}}$ (not on $T$ or $T_{0}$). Hence, given an error tolerance $\epsilon>0$, we can find $T^{{}^{\prime}}$ such that $g(T)\leq\epsilon$ for all $T\geq T^{{}^{\prime}}$. This will ensure adaptation in the case where channel probabilities remain fixed for windows longer than $T^{{}^{\prime}}$.
@@ I-A Utility functions
Another choice is $\phi(\boldsymbol{x})=\sum_{i=1}^{n}\log(1+\beta x_{i})$, where $\beta$ are given nonnegative weights. The logarithmic term introduces a form of proportional fairness [29, 30]. See also discussion of different utility functions in [29, 30, 38, 4, 35]
@@ I-B Related work
Network scheduling in stochastic environments is a widely considered problem in the literature on communication networks. Some examples include scheduling in wireless networks [23, 28], computer networks [37], vehicular networks [14] and unmanned aerial vehicle networks [31]. These works consider different goals such as optimal scheduling for queue-stability [27], power minimization [18], utility maximization [43].
@@
Multiple access, where many users access a limited number of communication channels, is an important problem in network scheduling. Here, it is desirable for users to be scheduled to avoid collisions. Transmission failures occur when the receiver is unable to decode the packet transmissions. This can occur, for example, when a fixed transmission power is used, but channel conditions have random and unknown fluctuations so that the received signal strength is insufficient for decoding. Different links can have different properties (such as different geographic distances to the receiver), so they can also have different success probabilities. The naive approach of assigning the users to links with the least failure probabilities leads to unfairness since, in such a scenario, users with high transmission failure probabilities will never be assigned. A common approach to solve this problem is to maximize a utility function of the time-averaged success. This problem has been well studied in the full information scenario when either the fluctuating channel conditions are known before transmission (called opportunistic scheduling) or when channel success probabilities are known in advance. Opportunistic scheduling has been considered using utility functions [36], Lyapunov drift [47], Frank-Wolfe [33, 1, 41], primal-dual [22, 45], and drift-plus-penalty [40]. The case when success probabilities are known in advance can be solved offline as a convex optimization problem using the mirror descent technique [20].
The problem becomes challenging when the success probabilities are unknown and only bandit feedback on transmission successes is available. As a result, the problem must be addressed by combining ideas from utility optimization over time averages with multi-armed bandit learning. Existing work on bandits with vector rewards and concave utility functions can be adapted to the single-channel case ($m=1$) of our problem (see [2, 19, 3]). However, these approaches have two main drawbacks. First, they do not consider the matching constraints present in our setting. Second, they rely on upper confidence bound (UCB) techniques and are not adaptive. In particular, these algorithms cannot adapt to changes in transmission failure probabilities without being explicitly informed of such changes. Hence, the goal of this paper is to develop adaptive algorithms for utility optimization with matching constraints under bandit feedback. This is achieved by combining the EXP3.S algorithm [7], which enables adaptive bandit learning, with Lyapunov optimization [42] to address the utility maximization objective. Although EXP3.S was originally developed for adversarial bandits, it has also been successfully applied in nonstationary stochastic environments [27].
@@
Existing work on using learning for scheduling problems in communications includes spectrum sharing [11, 21], network utility maximization with unknown utility functions under full feedback [25], and queue stability [27]. The literature on bandit learning-based scheduling primarily focuses on maximizing throughput [46]. The work of [25] considers network utility maximization where the utility functions are learned online, and [44] studies online resource allocation with the objective of maximizing $\alpha$-fairness. Neither of these works considers a bandit reward structure or matching-type constraints, and both differ from our problem formulation. The work of [10] studies an online optimization problem on transport polytopes. While their framework can be adapted to the Birkhoff polytope to incorporate matching constraints, it does not extend to bandit feedback settings. Moreover, none of the above works explicitly addresses the adaptiveness of the proposed algorithms.
@@
It should be noted that the concept of adaptiveness considered in this paper is different from the adversarial setting [8, 48, 7], although they have similarities. In adversarial settings, the guarantee of convergence is defined with respect to the rewards received throughout the time horizon. This is useful when no stochastic assumptions can be placed on the rewards. On the other hand, the adaptiveness considered in this paper is useful when we can place stochastic assumptions on the rewards, but the distributions may change from time to time. In this case, we can define an optimal strategy for each time frame in which the reward distribution remained constant. Hence, the goal in this frame is to learn the aforementioned strategy irrespective of the reward distributions of the past. An approach commonly used in problems of this flavor is minimizing dynamic regret [9, 16, 17]. Here, the regret is modified to account for the changing environments and the regret bounds are in terms of some measure that captures the degree of change. Various algorithms are developed for the setting with linear utility functions using optimizing in phases/episodes [9, 16], and sliding window-based algorithms [26, 17]. We utilize a simpler notion of adaptiveness and develop an algorithm for the case with general utility functions.
Multi-armed bandit (MAB) learning [34, 6] provides a principled framework for sequential decision-making under uncertainty. Classical stochastic bandit models focus on learning unknown mean rewards, often using upper confidence bound (UCB)–type algorithms [13, lattimore_szepesvári_2020]. In contrast, adversarial bandit models make no stochastic assumptions on rewards and are commonly addressed using randomized algorithms such as EXP3 [7]. Beyond these settings, numerous extensions have been studied, including linear, contextual, and combinatorial bandits [13].
Table I positions our work relative to the most closely related literature. To the best of our knowledge, no prior work simultaneously addresses utility optimization under bandit feedback with matching constraints while providing interval-based adaptiveness guarantees. Unlike classical bandit algorithms, which maximize scalar time-averaged reward over independent arms, our setting involves combinatorial matching constraints and optimization of a concave utility function of the time-average throughput vector. These structural features introduce nonlinear coupling across users and channels and prevent a direct application of standard bandit techniques. Furthermore, achieving adaptiveness under changing channel statistics requires incorporating EXP3.S-style exploration mechanisms within a Lyapunov optimization framework.
TABLE I: Comparison of work on Network Utility Optimization and Adaptive Bandits.
Work Utility Feedback Environment Matching Adaptive Bound
Constraints
Adaptive MAC General Lipschitz Bandit Stochastic Yes Yes $\mathcal{O}\left(\sqrt{\frac{\log(T)}{T}}\right)$
(This Paper)
Adaptive
MAC.CF General Lipschitz Bandit Stochastic Yes Yes $\mathcal{O}\left(\sqrt[3]{\frac{\log(T)}{T}}\right)$
(This Paper)
UCB MAC General Lipschitz Bandit Stochastic Yes No $\mathcal{O}\left(\sqrt{\frac{\log(T)}{T}}\right)$
(This Paper)
Neely [40] General Lipschitz Full Stochastic No Yes $\mathcal{O}\left(\sqrt{\frac{1}{T}}\right)$
Agarwal et al [2] General Lipschitz Bandit Stochastic No No $\mathcal{O}\left(\sqrt{\frac{\log(T)}{T}}\right)$
Neely [41] General Smooth Full Stochastic No No $\mathcal{O}\left(\frac{\log(T)}{T}\right)$
Garivier et al [26] Linear Bandit Stochastic No Yes $\mathcal{O}\left(\sqrt{\frac{\log(T)}{T}}\right)$
Ballu et al [10] General Lipschitz Full Stochastic Yes No $\mathcal{O}\left(\sqrt{\frac{1}{T}}\right)$
@@ I-C Our Contributions
  1. 1.
    Novel Problem Formulation: We develop and analyze algorithms to solve the problem of automatic link selection in multi-channel multiple access by combining ideas from multi-armed bandit learning and Lyapunov optimization. Although the classical multi-armed bandit (MAB) problem has been widely studied for maximizing linear utilities, such approaches can suffer from a lack of fairness in assignments when applied to the problem considered here. In contrast, our framework allows for smooth or nonsmooth concave utility functions.
  2. 2.
    Adaptive Algorithm with Performance Guarantees: We develop two adaptive algorithms that jointly address bandit learning, matching constraints, and utility maximization. The algorithms combine the EXP3.S algorithm for adversarial bandits [7] with Lyapunov optimization and techniques for online matching problems developed in [15]. We show that the first algorithm (Adaptive MAC) algorithm achieves an $\mathcal{O}(\sqrt{\log(T)/T})$ performance gap over any interval of $T$ consecutive time slots during which the (unknown) success probabilities remain fixed. This regret bound matches, up to logarithmic factors, the minimax lower bound for non-stationary multi-armed bandits established in [26]. Moreover, if the success probabilities change at some time $T_{0}$, the performance guarantees over the interval $\{T_{0},T_{0}+1,\dots,T_{0}+T-1\}$ are independent of the system behavior prior to $T_{0}$, even though the algorithm does not know the change point $T_{0}$. Hence, the algorithm is adaptive. Adaptive MAC achieves fast convergence but requires solving a convex optimization problem in each iteration, which is computationally expensive. To reduce computational complexity, we propose a second adaptive algorithm (Adaptive MAC.CF) whose iterations admit closed-form solutions. This approach trades off a slower convergence rate for significantly lower per-slot computational cost.
  3. 3.
    Special Cases: In the single-channel special case, we recover both fast convergence and closed form solutions for iterations via a simpler adaptive algorithm. When adaptation is not required, we also consider a UCB-based algorithm that achieves the same asymptotic convergence rate while relying only on simple max-weight-type decisions. The UCB algorithm highlights the advantages of adaptive algorithms in non-stationary environments. We further study the single-channel case with utility function $\phi(\boldsymbol{x})=min\{x_{1},\dots,x_{n}\}$, where a simple estimation-free mechanism achieves both convergence and adaptation.
  4. 4.
    Empirical evaluation: Through simulations, we demonstrate the fast convergence of the proposed adaptive algorithms and their ability to quickly adjust to changes in link success probabilities. In the conducted simulations, the Adaptive MAC.CF algorithm reduces the per-iteration computational complexity by approximately 36% compared to the Adaptive MAC algorithm. We also empirically compare our methods with prior work in the single-channel setting.